in triangle JKL, J is a right angle. point M is on Lk such that JM is perpendicular to LK. If LM = 6 and LK = 15, find JL. (simplify radicals.)
Answer by rothauserc(4718)

1) JL^2 + JK^2 = 225
:
2) JM^2 + 36 = JL^2
:
3) JM^2 + 81 = JK^2
:
By equation 1) JK^2 = 225 - JL^2
:
substitute for JK^2 in equation 3)
:
4) JM^2 + 81 = 225 - JL^2
:
By equation 2 JM^2 = JL^2 - 36
:
substitute for JM^2 in equation 4)
:
JL^2 - 36 + 81 = 225 - JL^2
:
2JL^2 = 180
:
JL^2 = 90
:
*****************************
JL = sqrt(90) = 3 * sqrt(10)
*****************************
: